Grade 12th passDiscuss with colleagues and IITians

a carnot engine absorbs 200J of heat from a reservoir at the temperature of normal boiling point of water and rejects heat to a reservoir at the temperature of the triple point off water. Find the heat rejected, the work done by the engine and the thermal efficiency.

MANOJ CHAUHAN
8 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
Q1 = 1000 J  taking from source
Q2 = 600 J   to sink
 
W = useful work done per cycle = Q1 - Q2 = 400 J
η = efficiency of the Carnot engine = W / Q1 =(Q1 - Q2)/ Q1 *100 = 40%
 
T1 = 127° C = 400°K
 
Q1 / Q2 =  T1/T2     =>  1000/600 = 400/T2
T2 = 240°K  =  - 33 °C
 
Alternately,   η = 0.4 = (T1 - T2)/T1
          So T1 - T2 = 400*0.4 = 160 K
                T2 = 240 K